### v1

# First-Hitting Location Laws as Boundary Observables of Drift–Diffusion Processes

Abstract
We investigate first-hitting location (FHL) statistics induced by drift–diffusion processes in domains with absorbing boundaries, and examine how such boundary laws give rise to intrinsic information observables. Rather than introducing explicit encoding or decoding mechanisms, information is viewed as emerging directly from the geometry and dynamics of stochastic transport through first-passage events. Treating the FHL as the primary observable, we characterize how geometry and drift jointly shape the induced boundary measure. In diffusion-dominated regimes, the exit law exhibits scale-free, heavy-tailed spatial fluctuations along the boundary, whereas a nonzero drift component introduces an intrinsic length scale that suppresses these tails and reorganizes the exit statistics. Within a generator-based formulation, the FHL arises naturally as a boundary measure induced by an elliptic operator, allowing explicit boundary kernels to be derived analytically in canonical geometries. Planar absorbing boundaries in different ambient dimensions are examined as benchmark cases, illustrating how directed transport regularizes diffusion-driven fluctuations and induces qualitative transitions in boundary statistics. Overall, the present work provides a unified structural framework for first-hitting location laws and highlights FHL statistics as natural probes of geometry, drift, and irreversibility in stochastic transport.
I Introduction
First-passage phenomena play a central role in nonequilibrium statistical physics, stochastic transport, and reaction–diffusion theory [21, 3, 5, 22, 4]. Classical questions concern the time at which a diffusing particle first reaches a target, the probability of absorption, or the survival probability up to a given time. These quantities have been extensively studied for Brownian motion and its generalizations, with applications ranging from chemical kinetics and biophysics to search processes and transport in disordered media.
In many physical settings, however, a boundary event is not fully characterized by its occurrence time alone. When absorption takes place on an extended boundary, the location at which the particle first hits the boundary constitutes an additional observable that encodes both geometric and dynamical information. This first-hitting location (FHL) observable is particularly relevant in spatially structured environments—such as interfaces, membranes, or planar detectors—where spatial heterogeneity along the boundary cannot be neglected [11]. Despite its conceptual simplicity, the statistical characterization of first-hitting locations has received comparatively less attention than that of first-passage times; in the drift-free case, it is closely connected to harmonic measure and its geometric regularity [16].
From a probabilistic perspective, the distribution of the first-hitting location on an absorbing boundary is a classical object, closely related to harmonic measure and boundary fluxes induced by elliptic operators [7, 17, 23]. In the absence of drift, it coincides with the harmonic measure of the domain and admits a representation in terms of the Poisson kernel associated with the Laplacian. More generally, for diffusions with drift, the hitting distribution can be characterized as the boundary measure induced by the corresponding elliptic operator with Dirichlet boundary conditions, via standard Green-function representations [7, 17, 23]. These connections form part of the classical foundation of potential theory and provide the conceptual backdrop for the present work. What is new in the present work is not the identification of boundary hitting measures per se, but a structural representation of first-hitting location laws that makes explicit how geometry and drift enter boundary observables. This representation organizes boundary statistics through induced elliptic measures, clarifying the role of directed transport without relying on geometry-specific constructions or case-by-case calculations.
In much of the existing literature, however, boundary-hitting statistics are primarily analyzed through temporal observables, such as survival probabilities, first-passage times, or boundary fluxes [22, 21, 9]. Even in simple geometries where explicit solutions are available, the normalized distribution of the FHL is often left implicit, rather than treated as a primary observable in its own right [22, 21]. This imbalance becomes more pronounced in the presence of drift, where directed transport reorganizes diffusion-induced spatial fluctuations and reshapes the resulting boundary laws.
The present work is primarily conceptual and structural in nature. Rather than introducing new stopping-time problems or cataloging boundary-hitting distributions, we focus on the representation of first-hitting location statistics as boundary-induced observables arising from stochastic transport. From this viewpoint, geometry and drift enter directly at the level of the induced boundary measure, rather than only through auxiliary temporal statistics.
To this end, we adopt a generator-based formulation for drift–diffusion processes with absorbing boundaries. Within this framework, first-hitting location laws arise naturally as boundary measures induced by elliptic operators, obtained from the outward normal derivative of the associated Green function. The emphasis is on providing a unified and transparent representation of boundary location statistics, rather than on developing new boundary-hitting techniques.
Classical first-passage and boundary-hitting theories have long emphasized the distribution of stopping times. In contrast, our focus is on the spatial structure of the induced boundary measures themselves, which encode information about transport geometry, irreversibility, and boundary symmetry. This shift in emphasis is particularly natural in nonequilibrium settings, where drift breaks time-reversal symmetry and renders boundary observables sensitive probes of transport organization.
Conceptually, our model is intentionally minimal. We couple simple irreversible transport (drift-diffusion) with absorbing boundaries that admit closed-form boundary measures. This framework clarifies the interplay between geometry and transport without modeling detailed microscopic effects.
As concrete realizations of this framework [9], we derive explicit closed-form expressions for the first-hitting location density in canonical geometries, including two- and three-dimensional planar absorbing boundaries, allowing for arbitrary constant drift directions. These settings serve as analytically tractable representatives of distinct symmetry classes rather than as universal models for all geometries. In particular, the planar cases make transparent how directed transport suppresses diffusion-induced spatial fluctuations along the boundary and introduces a characteristic length scale $\ell_{u}=\sigma^{2}/\|\bm{v}\|=1/u$, which separates diffusion-dominated, scale-free behavior from drift-regularized regimes with exponentially screened tails.
Scope and limitations.
The present work focuses on planar absorbing geometries as canonical benchmarks for boundary-induced first-hitting location statistics. While the generator-based formulation and the qualitative transition between diffusion-dominated and drift-regularized regimes are not restricted to planar domains, explicit closed-form boundary kernels are derived here only for the half-space geometry. Extensions to curved boundaries, time-dependent drift fields, and fully information-theoretic performance measures such as channel capacity are beyond the scope of this study and will be addressed elsewhere. Our aim is instead to isolate the minimal structural mechanisms by which geometry and drift jointly shape boundary exit laws and their induced observables.
The remainder of the paper is organized as follows. Section II introduces the stochastic transport model and defines the first-hitting location observables considered in this work. Section III develops a generator–Green-function framework that provides the analytical foundation for characterizing boundary exit laws via elliptic boundary-value problems. In Sec. IV, explicit boundary kernels are derived for planar absorbing geometries in two and three dimensions, serving as canonical benchmarks for boundary-induced exit statistics. Section V examines the asymptotic structure of these exit laws and the information observables they induce, and Sec. VI concludes the paper.
Table 1: Parabolic PDE versus generator-based routes to first-hitting location laws
Parabolic PDE route Generator–Green-function route
Step 1 Construct a free-space fundamental solution for a parabolic partial differential equation (PDE). Pass to the infinitesimal generator of the Markov semigroup.
Step 2 Impose absorbing boundary conditions (e.g., by image methods). Solve an elliptic boundary-value problem with Dirichlet data.
Step 3 Compute boundary flux and integrate over all time. Extract the boundary kernel from the normal derivative of the associated Green function.
II Model
The formulation adopted here is closely related to generator-based descriptions of stochastic transport with absorbing boundaries [7, 9]. Related approaches have also appeared in the molecular communication literature [14]. Here we intentionally adopt a more abstract, geometry-forward perspective. Rather than emphasizing system-level communication performance, our focus is on the structural role of geometry and drift in shaping boundary-induced statistics of stochastic transport.
Notation and dimensional conventions.
Throughout the paper, $d$ denotes the spatial dimension; its precise interpretation is clear from context. We work in an ambient space $\mathbb{R}^{d}$ with $d\geq 2$. Particle trajectories evolve in a domain $\Omega\subset\mathbb{R}^{d}$ with absorbing boundary $\partial\Omega$. The boundary is therefore a (d1)(d-1)-dimensional manifold, and boundary-parallel coordinates are denoted by $\bm{r}\in\mathbb{R}^{d-1}$. A summary of notation is provided in Table II for reference.
The diffusion coefficient is denoted by $D>0$, with the standard parametrization $D=\sigma^{2}/2$, following the conventional usage in statistical physics whereby $D$ characterizes the macroscopic spreading rate of the process and appears as the prefactor of the Laplacian in the diffusion equation. In the equivalent stochastic differential equation formulation, the microscopic noise amplitude enters via $\mathrm{d}X_{t}=\sigma\,\mathrm{d}B_{t}$, and the two descriptions are related precisely through $D=\sigma^{2}/2$. Accordingly, throughout this work we reserve the term diffusion coefficient for $D$, while referring to $\sigma$ as the noise strength. For later use, it is convenient to introduce the dimensionless drift parameter
$\bm{u}:=\bm{v}/\sigma^{2},\quad\bm{u}=(\bm{u}_{\parallel},u_{d}),$ (1)
which naturally appears in the boundary kernel. With this convention, the cases $d=2$ and $d=3$ correspond respectively to an absorbing line and an absorbing plane, while higher-dimensional extensions fit within the same unified framework.
II.1 Stochastic transport model
Let $\{\mathbf{X}_{t}\}_{t\geq 0}$ be an Itô diffusion [18] in $\Omega$ with absorbing boundary $\partial\Omega$. The process starts from $\mathbf{X}_{0}=\bm{x}_{0}\in\Omega$ and satisfies the evolution equation:
$\mathrm{d}\mathbf{X}_{t}=\bm{v}\,\mathrm{d}t+\sigma\,\mathrm{d}\mathbf{B}_{t},\quad t\geq 0,$ (2)
where $\mathbf{B}_{t}$ is a $d$-dimensional standard Brownian motion. In the representative geometries studied in Sec. IV, we consider constant drift $\bm{v}$ and isotropic diffusion with noise strength $\sigma$.
At the level of densities, the transition probability density $p(\bm{x},t\mid\bm{x}_{0})$ of the killed process in $\Omega$ satisfies the advection–diffusion equation with absorbing boundary condition $p=0$ on $\partial\Omega$ [22]:
$\partial_{t}p=-\nabla\!\cdot(\bm{v}\,p)+\frac{\sigma^{2}}{2}\,\Delta p,\quad\mathbf{x}\in\Omega.$ (3)
II.2 First-hitting observables
Define the exit time from the domain $\Omega$ by
$\tau_{\Omega}:=\inf\{t>0:\mathbf{X}_{t}\in\partial\Omega\},$ (4)
and the corresponding boundary exit location by
$\mathbf{X}_{\tau_{\Omega}}\in\partial\Omega.$ (5)
The primary object of interest in this work is the first-hitting location, viewed as a boundary-induced observable. While boundary-hitting phenomena are often characterized through temporal quantities such as survival probabilities or first-passage times, we instead treat the normalized distribution of the boundary exit location as the central observable [21, 4, 15] . From this perspective, geometry and drift act directly on the induced boundary measure, rather than entering only through auxiliary stopping-time statistics.
Let $\omega^{\bm{x}}$ denote the exit measure on $\partial\Omega$ associated with a trajectory starting from $\bm{x}$. When $\partial\Omega$ is sufficiently regular, $\omega^{\bm{x}}$ is absolutely continuous with respect to the surface measure $\mathrm{d}S_{\bm{y}}$ on $\partial\Omega$ [7, 19, 16] , and we write
$\omega^{\bm{x}}(\mathrm{d}\bm{y})=\mathcal{K}(\bm{x},\bm{y})\,\mathrm{d}S_{\bm{y}},\quad\bm{y}\in\partial\Omega,$ (6)
where $\mathcal{K}(\bm{x},\bm{y})$ is the associated boundary kernel, denoted throughout by $\mathcal{K}(\bm{x},\bm{y})$ to avoid confusion with the modified Bessel function $K_{\nu}(\cdot)$.
Equivalently, for any test function $g$ defined on $\partial\Omega$, we have
$\mathbb{E}_{\bm{x}}\!\left[g(\mathbf{X}_{\tau_{\Omega}})\right]=\int_{\partial\Omega}g(\bm{y})\,\mathcal{K}(\bm{x},\bm{y})\,\mathrm{d}S_{\bm{y}}.$ (7)
The main technical objective of this paper is to determine the boundary kernel $\mathcal{K}(\bm{x},\bm{y})$ explicitly in representative geometries and to relate it to elliptic boundary-value problems [12] driven by the generator of the diffusion.
At this stage, the exit time $\tau_{\Omega}$ is treated as an internal variable and is marginalized out in the analysis, unless otherwise stated.
III A unified framework for determining first-hitting location laws
This section presents a unified route, summarized in Table I, for determining FHL laws induced by drift–diffusion processes in domains with absorbing boundaries. Related generator-based descriptions of boundary observables and boundary local time have been extensively studied in the context of reflected and absorbed diffusions; see, e.g., [10].
We refer to the exit location on the boundary as the FHL. The approach relies on standard tools from stochastic analysis and potential theory, in particular, the infinitesimal generator of the diffusion semigroup and Dynkin’s formula [7, 19] . By reformulating the problem at the level of the generator, the determination of FHL laws is reduced to properties of an elliptic boundary-value problem.
At a conceptual level, the procedure consists of three steps, summarized in Table 1. The essential observation is that once the elliptic Dirichlet Green function $\mathcal{G}(\bm{x},\bm{y})$ associated with the generator is available, the exit law on the boundary admits a kernel representation of the form [7, 19, 9]:
$\mathcal{K}(\bm{x},\bm{y})=-\,\partial_{n(\bm{y})}\mathcal{G}(\bm{x},\bm{y}),\quad\bm{y}\in\partial\Omega,$ (8)
where $\partial_{n(\bm{y})}$ denotes differentiation along the outward normal at the boundary point $\bm{y}$.
In contrast to time-dependent (parabolic) approaches [21, 22] that analyze the evolution equation directly, the generator-based formulation removes the time variable from the outset. As a result, the FHL law is obtained without performing any integration over time.
We now derive the generator-based representation and specify the assumptions under which it holds.
III.1 Infinitesimal generator and Markov semigroup
Let $\{\mathbf{X}_{t}\}_{t\geq 0}$ be the Itô diffusion introduced in Sec. II, evolving in $\Omega$ with absorbing boundary $\partial\Omega$. Throughout this work, we restrict attention to the constant-drift, isotropic-diffusion setting specified in Sec. II.
For a time-homogeneous Markov process, the operators
$P_{t}f(\bm{x}):=\mathbb{E}_{\bm{x}}\!\left[f(\mathbf{X}_{t})\right],\quad t\geq 0,$ (9)
define a Markov semigroup $\{P_{t}\}_{t\geq 0}$ acting on suitable test functions $f$. The infinitesimal generator $\mathcal{L}$ is defined on its domain $\mathcal{D}(\mathcal{L})$ by
$\mathcal{L}f:=\lim_{t\to 0^{+}}\frac{P_{t}f-f}{t},\quad f\in\mathcal{D}(\mathcal{L}).$ (10)
Applying Itô’s formula to $f(\mathbf{X}_{t})$ for $f\in C^{2}$ and using the stochastic differential equation $\mathrm{d}\mathbf{X}_{t}=\bm{v}\,\mathrm{d}t+\sigma\,\mathrm{d}\mathbf{B}_{t}$, we obtain
$\mathrm{d}f(\mathbf{X}_{t})=\$$(\nabla f)^{\mathsf{T}}\mathrm{d}\mathbf{X}_{t}+\frac{1}{2}(\mathrm{d}\mathbf{X}_{t})^{\mathsf{T}}\mathrm{Hess}(f)(\mathrm{d}\mathbf{X}_{t})$ (11)
$=\$$\Big[\bm{v}^{\mathsf{T}}\nabla f+\frac{\sigma^{2}}{2}\,\Delta f\Big]\mathrm{d}t+(\nabla f)^{\mathsf{T}}\sigma\,\mathrm{d}\mathbf{B}_{t}.$ (12)
Taking expectations and using $\mathbb{E}_{\bm{x}}[\mathrm{d}\mathbf{B}_{t}]=0$, we arrive at the explicit form of the generator
$\mathcal{L}f(\bm{x})=\bm{v}^{\mathsf{T}}\nabla f(\bm{x})+\frac{\sigma^{2}}{2}\,\Delta f(\bm{x}),$ (13)
which is a second-order elliptic operator with constant coefficients.
We next consider the Dirichlet boundary-value problem
$\left\{\begin{array}[]{ll}\mathcal{L}\phi=0&\text{in }\Omega,\\ \phi=g&\text{on }\partial\Omega,\end{array}\right$. (14)
where $g$ is a prescribed boundary function on $\partial\Omega$.
III.2 Green-function representation of the exit measure
From the definition of the generator,
$\mathbb{E}_{\bm{x}}[f(\mathbf{X}_{t})]=f(\bm{x})+\mathbb{E}_{\bm{x}}\!\left[\int_{0}^{t}\mathcal{L}f(\mathbf{X}_{s})\,\mathrm{d}s\right],$ (15)
for deterministic time $t\in\mathbb{R}^{+}$. Dynkin’s formula extends this identity to stopping times. Let $\tau_{\Omega}$ denote the exit time from $\Omega$, and assume $\mathbb{E}_{\bm{x}}[\tau_{\Omega}]<\infty$. Then for $f\in C^{2}$,
$\mathbb{E}_{\bm{x}}[f(\mathbf{X}_{\tau_{\Omega}})]=f(\bm{x})+\mathbb{E}_{\bm{x}}\!\left[\int_{0}^{\tau_{\Omega}}\mathcal{L}f(\mathbf{X}_{s})\,\mathrm{d}s\right].$ (16)
If $\phi$ solves the Dirichlet problem (14), then $\mathcal{L}\phi=0$ in $\Omega$. Setting $f=\phi$ in (16) yields
$\phi(\bm{x})=\mathbb{E}_{\bm{x}}[g(\mathbf{X}_{\tau_{\Omega}})].$ (17)
Introducing the exit measure $\omega^{\bm{x}}$ on $\partial\Omega$, this relation can be written as
$\phi(\bm{x})=\int_{\partial\Omega}g(\bm{y})\,\omega^{\bm{x}}(\mathrm{d}\bm{y}).$ (18)
On the other hand, solutions of elliptic boundary-value problems admit Green function representations. Let $\mathcal{G}(\bm{x},\bm{y})$ denote the Dirichlet Green function of the elliptic operator $\mathcal{L}$. Then
$\phi(\bm{x})=\int_{\partial\Omega}g(\bm{y})\,\bigl[-\partial_{n(\bm{y})}\mathcal{G}(\bm{x},\bm{y})\bigr]\,\mathrm{d}S_{\bm{y}}.$ (19)
Comparing (18) and (19) yields the representation
$\omega^{\bm{x}}(\mathrm{d}\bm{y})=\mathcal{K}(\bm{x},\bm{y})\,\mathrm{d}S_{\bm{y}},\quad\mathcal{K}(\bm{x},\bm{y})=-\,\partial_{n(\bm{y})}\mathcal{G}(\bm{x},\bm{y}),$ (20)
which expresses the exit law on $\partial\Omega$ as a boundary kernel given by the outward normal derivative of the elliptic Green function. Explicit evaluations of $\mathcal{K}(\bm{x},\bm{y})$ for representative geometries are given in Sec. IV.
IV Calculating boundary kernels in representative geometries
[…]
Figure 1: Schematic illustration of a FHL process in an absorbing half-space. A drift–diffusion trajectory is initialized at a normal distance $\lambda$ from the absorbing boundary and evolves in the bulk until it is terminated upon first contact with the boundary. The spatial location of this exit event defines the primary observable of interest. The figure provides an interpretive visualization of how stochastic transport maps initial conditions to boundary exit measures.
This section applies the general exit-law representation (20) to geometries in which the boundary kernel $\mathcal{K}(\bm{x},\bm{y})$ can be evaluated in closed form [21, 7, 9]. We focus on the absorbing half-space, which provides the simplest extended boundary geometry and serves as a canonical local model for smooth interfaces in the limit of vanishing curvature.
Within this setting, two cases are of primary interest: (i) a two-dimensional system with an absorbing line (Fig. 1) and (ii) a three-dimensional system with an absorbing plane (Fig. 2). Both cases admit explicit boundary kernels under constant drift and isotropic diffusion, yielding analytically tractable expressions that make clear how directed transport regularizes diffusion-induced spatial fluctuations. These hyper-planar results constitute one of the main explicit contributions of the present work and serve as benchmark solutions for drift–diffusion exit statistics.
[…]
Figure 2: Schematic illustration of a drift–diffusion process in a half-space with an absorbing boundary (shown for $d=3$). A trajectory initialized at $\bm{x}_{0}\in\Omega$ evolves in the bulk under a constant drift $\bm{v}$ and isotropic diffusion, and terminates upon first reaching the absorbing boundary $\partial\Omega$ located at a normal distance $\lambda$. The corresponding exit time $\tau_{\Omega}$ and boundary location $\mathbf{X}_{\tau_{\Omega}}$ define the induced boundary law. The figure highlights the geometric setting and variables underlying the first-hitting location framework analyzed in the text.
IV.1 Absorbing half-space: reduction to an elliptic problem
We specialize the stochastic transport model defined in Eq. (2) to the case of an absorbing half-space $\Omega$ with boundary $\partial\Omega$. Under the constant-drift and isotropic-diffusion assumptions of Eq. (3), the infinitesimal generator $\mathcal{L}$ of the process is a second-order elliptic operator with drift:
$\mathcal{L}=\bm{v}^{\mathsf{T}}\nabla+\frac{\sigma^{2}}{2}\Delta.$ (21)
The boundary kernel $\mathcal{K}(\bm{x},\bm{y})$ defined in Eq. (6) is obtained from the outward normal derivative of the Green function associated with the Dirichlet problem $\mathcal{L}\phi=0$ in $\Omega$.
To determine $\mathcal{K}(\bm{x},\bm{y})$, we reduce the problem to a Helmholtz form by exploiting the dimensionless drift $\bm{u}$ introduced in Sec. II. Applying the exponential change of variables $\psi(\bm{x})=\exp(\bm{u}^{\mathsf{T}}\bm{x})\phi(\bm{x})$, the Dirichlet problem transforms into a Helmholtz-type equation for $\psi$:
$\left\{\begin{array}[]{ll}(\Delta-\|\bm{u}\|^{2})\psi=0&\text{in }\Omega,\\ \psi=\tilde{g}&\text{on }\partial\Omega,\end{array}\right$. (22)
where $\tilde{g}(\bm{y})=e^{\bm{u}^{\mathsf{T}}\bm{y}}g(\bm{y})$ [22, 7, 19]. By adopting the normalized parameter $\bm{u}$, the diffusion strength $\sigma^{2}$ is implicitly absorbed into the geometric scale. Let $\tilde{\mathcal{G}}(\bm{x},\bm{y})$ denote the Green function associated with (22). The boundary kernel $\mathcal{K}(\bm{x},\bm{y})$ then admits the following exact representation:
$\mathcal{K}(\bm{x},\bm{y})=e^{\bm{u}^{\mathsf{T}}(\bm{y}-\bm{x})}\bigl[-\partial_{n(\bm{y})}\tilde{\mathcal{G}}(\bm{x},\bm{y})\bigr],\quad\bm{y}\in\partial\Omega.$ (23)
Figure 3: Boundary hitting distributions in the drift-free regime ($\bm{u}=0$), shown for two- and three-dimensional boundaries. In the absence of drift, the first-hitting location law is isotropic and exhibits scale-free algebraic decay, corresponding to a Cauchy-type boundary measure [7] without an intrinsic length scale.
IV.2 Two dimensional case: absorbing line
For $d=2$, the Helmholtz Green function in the half-space admits the image representation [20]:
$\tilde{\mathcal{G}}(\bm{x},\bm{y})=\frac{1}{2\pi}\!\left[K_{0}(\|\bm{u}\|\mathcal{R}_{1})-K_{0}(\|\bm{u}\|\mathcal{R}_{2})\right],$ (24)
with
$\mathcal{R}_{1}$ $=\sqrt{(x_{1}-\xi)^{2}+(x_{2}-\eta)^{2}},$ (25a)
$\mathcal{R}_{2}$ $=\sqrt{(x_{1}-\xi)^{2}+(x_{2}+\eta)^{2}},$ (25b)
where $\mathcal{R}_{2}$ corresponds to the standard image point associated with the absorbing boundary.
Fix $\bm{x}=(x_{1},0)$ and $\bm{y}=(\xi,\lambda)$, where $\lambda>0$ denotes the normal distance from the initial point to the boundary. Differentiation in the outward normal direction yields
$\partial_{n(\bm{y})}\tilde{\mathcal{G}}(\bm{x},\bm{y})=\frac{\|\bm{u}\|\lambda}{\pi}\frac{K_{1}(\|\bm{u}\|\rho)}{\rho},\penalty 10000\ \rho=\sqrt{(x_{1}-\xi)^{2}+\lambda^{2}}.$ (26)
Substituting this expression into the exponential tilting representation (23), the boundary kernel along the absorbing line is given explicitly by
$\mathcal{K}(\bm{x},\bm{y})=\frac{\|\bm{u}\|\lambda}{\pi}\exp\!\bigl\{\bm{u}^{\mathsf{T}}(\bm{y}-\bm{x})\bigr\}\frac{K_{1}(\|\bm{u}\|\rho)}{\rho},\quad\bm{y}\in\partial\Omega.$ (27)
IV.3 Three dimensional case: absorbing plane
For $d=3$, the Helmholtz Green function in the half-space is given by [20]:
$\tilde{\mathcal{G}}(\bm{x},\bm{y})=\frac{e^{-\|\bm{u}\|\mathcal{R}_{1}}}{4\pi\mathcal{R}_{1}}-\frac{e^{-\|\bm{u}\|\mathcal{R}_{2}}}{4\pi\mathcal{R}_{2}},$ (28)
with
$\mathcal{R}_{1}$ $=\sqrt{(x_{1}-\xi)^{2}+(x_{2}-\eta)^{2}+(x_{3}-\zeta)^{2}},$ (29a)
$\mathcal{R}_{2}$ $=\sqrt{(x_{1}-\xi)^{2}+(x_{2}-\eta)^{2}+(x_{3}+\zeta)^{2}},$ (29b)
where $\mathcal{R}_{2}$ corresponds to the standard image construction associated with the absorbing boundary.
Fix $\bm{x}=(x_{1},x_{2},0)$ and $\bm{y}=(\xi,\eta,\lambda)$, where $\lambda>0$ denotes the normal distance from the initial point to the boundary. After taking the outward normal derivative and applying the exponential tilting relation (23), the boundary kernel on the absorbing plane takes the explicit form
$\mathcal{K}(\bm{x},\bm{y})=\frac{\lambda}{2\pi}\exp\!\bigl\{\bm{u}^{\mathsf{T}}(\bm{y}-\bm{x})\bigr\}e^{-\|\bm{u}\|\rho}\frac{1+\|\bm{u}\|\rho}{\rho^{3}},$ (30)
where $\rho=\|\bm{y}-\bm{x}\|$.
Figure 4: Boundary hitting distributions in the presence of drift ($\bm{u}\neq 0$), shown for two- and three-dimensional boundaries. Drift introduces an intrinsic length scale and exponentially suppresses large boundary displacements, leading to localized and anisotropic hitting profiles consistent with the drift-regularized analytical kernels.
IV.4 General $(d+1)$-dimensional structure
The explicit results in two and three dimensions reveal a common, dimension-independent structure of boundary exit laws in the absorbing half-space. In both cases, the ambient space has dimension $d+1$, while the induced boundary kernel lives on a $d$-dimensional absorbing interface. Guided by these expressions, we are naturally led to the conjecture that in $\mathbb{R}^{d+1}$ the boundary kernel admits the universal form
$\mathcal{K}^{(d)}(\bm{r};\bm{u},\lambda)$ $=2\lambda\frac{\|\bm{u}\|^{\frac{d+1}{2}}}{(2\pi)^{\frac{d+1}{2}}}\exp\!\big(\bm{u}_{\parallel}^{\mathsf{T}}\bm{r}-u_{d}\lambda\big)$
$\quad\times\frac{K_{\frac{d+1}{2}}\!\left(\|\bm{u}\|\sqrt{\|\bm{r}\|^{2}+\lambda^{2}}\right)}{\left(\sqrt{\|\bm{r}\|^{2}+\lambda^{2}}\right)^{\frac{d+1}{2}}}.$ (31)
Here $\bm{r}\in\mathbb{R}^{d}$ denotes the tangential displacement along the absorbing boundary.
While (IV.4) is stated as a conjecture, its structure is strongly constrained by symmetry and scaling considerations. In particular, translational invariance along the absorbing boundary, rotational symmetry in the $d$ boundary-parallel directions, and the exponential tilting induced by the drift together single out the modified Bessel kernel appearing above. From this perspective, (IV.4) should be viewed as the natural extension of the exact two- and three-dimensional results to general boundary dimension, rather than as a heuristic guess.
As a sanity check, we specialize the conjectured expression (IV.4) to the case $d=1$, corresponding to a two-dimensional ambient space with a one-dimensional absorbing boundary. In this setting, the tangential displacement reduces to a scalar $r\in\mathbb{R}$, and the modified Bessel function appearing in (IV.4) becomes $K_{1}$ [1, 19] . Substituting $d=1$ and simplifying the prefactors, the general formula reduces exactly to
$\mathcal{K}^{(1)}(r;\bm{u},\lambda)=\frac{\|\bm{u}\|\lambda}{\pi}\exp\!\big(u_{\parallel}r-u_{d}\lambda\big)\frac{K_{1}\!\left(\|\bm{u}\|\sqrt{r^{2}+\lambda^{2}}\right)}{\sqrt{r^{2}+\lambda^{2}}},$
which coincides with the explicit boundary kernel obtained for the absorbing line in Sec. IV.2, cf. (27).
A further consistency check is obtained by specializing (IV.4) to the case $d=2$, corresponding to a three-dimensional ambient space with a two-dimensional absorbing boundary. In this case, the modified Bessel function reduces to the half-integer order $K_{3/2}$, which admits a closed-form representation. Substituting $d=2$ into (IV.4) and simplifying the resulting expression using the identity for $K_{3/2}$ yields
$\mathcal{K}^{(2)}(\bm{r};\bm{u},\lambda)=\frac{\lambda e^{\bm{u}_{\parallel}^{\mathsf{T}}\bm{r}-u_{d}\lambda}}{2\pi}\cdot\frac{e^{-\|\bm{u}\|\rho}(1+\|\bm{u}\|\rho)}{\rho^{3}},$ (32)
where $\rho=\sqrt{\|\bm{r}\|^{2}+\lambda^{2}}$ denotes the slant distance. The above form coincides exactly with the boundary kernel obtained for the absorbing plane in Sec. IV.3, cf. (30). Together with the $d=1$ case, this agreement confirms that the conjectured form (IV.4) correctly reproduces the known closed-form results in both two and three dimensions.
V Asymptotic structure and information observables of exit laws
The planar results of Sec. IV provide explicit exit laws on the absorbing boundary, arising solely from the interplay between stochastic dynamics and geometry. Recalling the kernel representation established in Sec. III, the boundary exit measure $\omega^{\bm{x}}(\mathrm{d}\bm{y})$ can be written as
$\omega^{\bm{x}}(\mathrm{d}\bm{y})=\mathcal{K}(\bm{x},\bm{y})\,\mathrm{d}S_{\bm{y}},\qquad\bm{y}\in\partial\Omega,$ (33)
where $\mathcal{K}(\bm{x},\bm{y})$ denotes the boundary kernel and $\mathrm{d}S_{\bm{y}}$ is the surface measure on $\partial\Omega$. These explicit kernel representations allow us to go beyond a purely structural description of exit laws and to examine both their asymptotic organization across diffusion-dominated and drift-regularized regimes, as well as the information-theoretic observables they naturally induce.
Although no explicit encoding or decoding protocols are introduced, the induced boundary law supports canonical information-theoretic functionals. A representative example is the differential entropy of the first-hitting location distribution, which quantifies the dispersive extent of boundary exit events. We therefore regard such quantities as information observables intrinsically associated with first-passage phenomena.
From a broader statistical-physics perspective, these observables may be viewed as geometric manifestations of nonequilibrium first-passage dynamics, rather than as quantities tied to externally imposed communication protocols [2, 8, 21, 4] .
V.1 Asymptotic structure of planar exit laws
The explicit (hyper)-planar kernels derived in Sec. IV reveal a robust asymptotic structure of boundary exit laws that is not tied to a specific spatial dimension. At large tangential displacements along the absorbing boundary, the qualitative behavior of the exit law is governed primarily by the presence or absence of an intrinsic length scale, rather than by microscopic details of the stochastic transport mechanism.
In the drift-free case, the exit law coincides with harmonic measure [7, 19, 16]. The induced boundary kernel is scale-free and exhibits algebraic decay at large distances. For a $d$-dimensional absorbing boundary, the asymptotic tail takes the form
$\mathcal{K}^{(d)}(\bm{r})\sim\|\bm{r}\|^{-(d+1)},\qquad\|\bm{r}\|\to\infty,\quad(u=0),$ (34)
where $\bm{r}\in\mathbb{R}^{d}$ denotes the boundary-parallel displacement and $u=\|\bm{u}\|$ is the magnitude of the dimensionless drift. This heavy-tailed behavior reflects the absence of any intrinsic spatial scale and is a generic feature of boundary laws induced by harmonic kernels [21, 4]. As a consequence, higher-order moments diverge, and moment-based or variance-controlled descriptions fail, reflecting the intrinsically scale-free nature of the exit statistics. This asymptotic form follows directly from the general kernel representation (IV.4): setting $\nu=(d+1)/2$ and $\rho=\sqrt{\|\bm{r}\|^{2}+\lambda^{2}}$, the modified Bessel function appearing in (IV.4) has argument $z=\|\bm{u}\|\,\rho$, so that in the zero-drift limit $\|\bm{u}\|\to 0$ one has $z\to 0^{+}$ and the small-argument expansion $K_{\nu}(z)\sim\tfrac{1}{2}\Gamma(\nu)(2/z)^{\nu}$ [1]. The prefactor $\|\bm{u}\|^{\nu}$ in (IV.4) cancels the divergence of $K_{\nu}(z)$, yielding
$\mathcal{K}^{(d)}(\bm{r};\bm{0},\lambda)\propto(\|\bm{r}\|^{2}+\lambda^{2})^{-\nu},$ (35)
and hence $\mathcal{K}^{(d)}(\bm{r};\bm{0},\lambda)\sim\|\bm{r}\|^{-2\nu}=\|\bm{r}\|^{-(d+1)}$ as $\|\bm{r}\|\to\infty$.
When a nonzero drift component toward the absorbing boundary is present, an intrinsic length scale $\ell_{u}\sim u^{-1}$ emerges and qualitatively alters the asymptotic behavior. In this drift-regularized regime, the boundary kernel acquires an exponential screening factor and obeys
$\mathcal{K}^{(d)}(\bm{r})\sim\exp\bigl(-u\,\|\bm{r}\|\bigr)\,\|\bm{r}\|^{-(d+2)/2},\quad\|\bm{r}\|\to\infty,\penalty 10000\ u\neq 0.$ (36)
This exponential screening effect follows from the large-argument behavior of the modified Bessel function appearing in the planar kernel representation: for $\|\bm{r}\|\to\infty$ at fixed $u>0$, the Bessel argument grows linearly with $\|\bm{r}\|$, and the leading asymptotic form $K_{\nu}(z)\sim\sqrt{\pi/(2z)}\,e^{-z}$ yields both the exponential cutoff and the reduced algebraic prefactor. Together with the drift-free case, these two regimes are governed respectively by the small- and large-argument limits of the same modified Bessel function; the relevant asymptotic expansions are summarized in Appendix C.
Remark (Characteristic length scale).
The exponential cutoff introduces an intrinsic characteristic length scale (CLS)
$\ell_{u}=\frac{\sigma^{2}}{\|\bm{v}\|}=\frac{1}{u},\quad u=\|\bm{u}\|,$ (37)
which suppresses long excursions parallel to the absorbing boundary and restores finiteness of a broad class of induced information observables. From this viewpoint, drift toward the boundary acts as a robust regularization mechanism that interpolates between scale-free and exponentially screened exit statistics across spatial dimensions [22, 21]. The characteristic scale $\ell_{u}\sim u^{-1}$ delineates two asymptotic regimes. For tangential displacements $\|\bm{r}\|\ll\ell_{u}$, the boundary kernel retains the scale-free structure inherited from harmonic measure, whereas for $\|\bm{r}\|\gg\ell_{u}$ the exponential factor dominates and suppresses long excursions. In this sense, $\ell_{u}$ acts as a crossover length separating diffusion-dominated and drift-regularized behavior at the level of boundary exit laws.
V.2 Differential entropy of the exit location
Let $\Xi$ denote the boundary exit location,
$\Xi:=\mathbf{X}_{\tau_{\Omega}}\in\partial\Omega,\quad\tau_{\Omega}:=\inf\{t>0:\mathbf{X}_{t}\in\partial\Omega\},$ (38)
under an initial condition $\mathbf{X}_{0}=\bm{x}$. Whenever the exit law $\omega^{\bm{x}}$ is absolutely continuous with respect to the surface measure on $\partial\Omega$, we write $\omega^{\bm{x}}(\mathrm{d}\bm{y})=\mathcal{K}(\bm{x},\bm{y})\mathrm{d}S_{\bm{y}}$ and define the (differential) entropy [6] of the exit location as
$H(\Xi\,|\,\bm{x})=-\int_{\partial\Omega}\mathcal{K}(\bm{x},\bm{y})\,\log\mathcal{K}(\bm{x},\bm{y})\,\mathrm{d}S_{\bm{y}}.$ (39)
This quantity characterizes the intrinsic spatial uncertainty of the exit event along the absorbing boundary and serves as a natural information observable induced by the boundary-hitting statistics.
In diffusion-dominated regimes, the exit law exhibits an algebraically decaying heavy tail along the boundary. In the drift-free limit, the kernel reduces to a Cauchy-type form with power-law decay, as in (34). Despite the absence of finite moments of order one or higher, such heavy-tailed distributions possess a well-defined differential entropy that depends only logarithmically on the scale parameter. The pathological feature relevant to information transmission is therefore not entropy divergence, but rather the lack of an intrinsic spatial scale and the resulting dominance of rare but arbitrarily large tangential displacements. In this regime, moment-based characterizations and Gaussian approximations break down, and information measures become highly sensitive to tail behavior.
The introduction of drift fundamentally alters this picture. As demonstrated by the explicit planar kernels in Sec. IV, a nonzero drift component directed toward the absorbing boundary induces an exponential suppression of large boundary displacements. This screening effect suggests a characteristic length scale $\ell_{u}$, as defined in (37), which governs the effective spatial extent of the exit distribution. For $\|\bm{v}\|>0$, the exit law is no longer scale-free, and the spatial uncertainty of the exit location is controlled by $\ell_{u}$. In particular, $\ell_{u}$ sets the effective spatial support over which the exit density contributes appreciably to entropy-like functionals, so that increasing drift strength progressively localizes the boundary distribution and reduces the contribution of rare, large displacements.
This regularization mechanism is already apparent at the level of the boundary kernel itself: the emergence of an intrinsic length scale suppresses long tangential excursions and converts an intrinsically scale-free exit observable into one governed by controlled, length-scale-dependent statistics.
A fully quantitative analysis of entropy variation or input–output information measures (e.g., the Shannon capacity [6]) would require specifying an input ensemble, and therefore lies beyond the scope of the present work.
VI Conclusion and discussion
This work examined boundary exit laws associated with the location of first-passage events in drift–diffusion processes on extended domains, with particular emphasis on their asymptotic structure and induced information observables. Within a generator–Green-function framework, we derived explicit boundary kernels for planar absorbing geometries and characterized the resulting exit measures across diffusion-dominated and drift-regularized regimes. Viewed from this perspective, the exit location, rather than the exit time alone, emerges as a sensitive observable through which geometry and directed transport are imprinted on boundary statistics. In the absence of drift, the exit law reduces to harmonic measure and exhibits algebraic, scale-free decay along the boundary, whereas a nonzero drift component toward the boundary induces exponential screening and introduces a characteristic length scale $\ell_{u}=\sigma^{2}/\|\bm{v}\|=1/u$. This transition is manifest directly at the level of the boundary kernel and persists across spatial dimensions.
Beyond the availability of explicit formulas, the analysis highlights a structural distinction intrinsic to first-passage exit statistics. When drift is absent, exit locations are governed solely by geometry and diffusive spreading, leading to scale-free behavior dominated by rare but long tangential excursions. The introduction of drift qualitatively alters this picture by generating an intrinsic length scale that suppresses such excursions and localizes the exit distribution. In this sense, first-hitting location statistics provide a concrete boundary manifestation of irreversibility in stochastic transport, reflecting how drift reorganizes spatial fluctuations at absorbing interfaces. Information-theoretic quantities, such as the differential entropy of the exit location, serve as convenient diagnostics of this transition, reflecting changes in boundary statistics induced jointly by stochastic transport and geometry.
The planar results presented here provide benchmark solutions that capture the leading-order structure of boundary exit laws in more general geometries, with curvature and global shape effects entering as higher-order corrections. More broadly, the framework developed in this work identifies first-hitting location statistics as a natural class of boundary-induced observables through which geometric and dynamical aspects of stochastic transport can be examined within a unified setting. Extensions to curved absorbing boundaries and to time-dependent drift constitute natural directions for future work, where additional geometric and dynamical scales are expected to further enrich the structure of boundary exit laws.
Acknowledgements.
This work was supported by the National Science and Technology Council of Taiwan under Grant No. NSTC 113-2115-M-008-013-MY3.
Data Availability
The data that support the findings of this study are available from the corresponding author upon reasonable request.
Author Contributions
Yen-Chi Lee conceptualized the research, developed the mathematical framework, derived the analytical kernels, and wrote the manuscript.
Appendix A Zero-drift limit and emergence of Cauchy exit laws
Throughout the main text, the exit law on $\partial\Omega$ is written as a boundary measure $\omega^{\bm{x}}(\mathrm{d}\bm{y})=\mathcal{K}(\bm{x},\bm{y})\,\mathrm{d}S_{\bm{y}}$. In the appendices we sometimes parameterize $\partial\Omega$ by local coordinates $y\in\mathbb{R}^{d-1}$ (e.g., $y=r$ for $d=2$ and $y=\bm{r}$ for $d=3$), in which case $\mathrm{d}S_{\bm{y}}=\mathrm{d}y$ and the same law is written in probability density function (PDF) form $f_{\Xi}(y)=\mathcal{K}(\bm{x},\bm{y}(y))$. More generally, for a coordinate map $y\mapsto\bm{y}(y)$, the density takes the form $f_{\Xi}(y)=\mathcal{K}(\bm{x},\bm{y}(y))\,J(y)$, where $J(y)$ denotes the surface Jacobian.
This appendix provides a technical derivation of the zero-drift limit ($v\to 0$) for the drift–diffusion FHL laws, establishing how the boundary exit distribution converges to a multidimensional Cauchy-type law characterized by algebraic, heavy-tailed decay [13]. This limiting behavior constitutes the mathematical foundation for the scale-free, diffusion-dominated regime analyzed in Secs. IV and V. We explicitly treat planar absorbing geometries in two and three spatial dimensions—parameterized by boundary coordinates $r\in\mathbb{R}$ and $\bm{r}\in\mathbb{R}^{2}$, respectively, with an initial normal distance $\lambda>0$. Notice that the plane-shape absorbing boundaries serve as canonical local models for general smooth boundaries and illustrate the fundamental mechanism by which heavy-tailed statistics emerge from first-passage processes.
A.1 Two-dimensional geometry: absorbing line
In two spatial dimensions, the absorbing boundary is a line. For a drift–diffusion process with constant drift $\bm{v}=(v_{1},v_{2})$, the boundary kernel associated with the exit law can be written in the form [14]
$\mathcal{K}^{(1)}(r;\bm{v})={}$ $\frac{\|\bm{v}\|\lambda}{\sigma^{2}\pi}\exp\!\left(-\frac{v_{2}\lambda}{\sigma^{2}}\right)\exp\!\left(-\frac{v_{1}r}{\sigma^{2}}\right)$ (40)
$\times\frac{K_{1}\!\left(\frac{\|\bm{v}\|}{\sigma^{2}}\sqrt{r^{2}+\lambda^{2}}\right)}{\sqrt{r^{2}+\lambda^{2}}},$
where $K_{1}(\cdot)$ denotes the modified Bessel function of the second kind. This kernel satisfies $\omega^{\bm{x}}(\mathrm{d}r)=\mathcal{K}^{(1)}(r;\bm{v})\,\mathrm{d}r$ for $\bm{x}=(0,\lambda)$.
To examine the zero-drift regime, we take the limit $\|\bm{v}\|\to 0$. Using the standard asymptotic relation [1]:
$\lim_{x\to 0}xK_{1}(x)=1,$ (41)
the exponential prefactors in (40) converge to unity, and the leading contribution arises solely from the Bessel kernel. A direct calculation yields
$\lim_{\|\bm{v}\|\to 0}\mathcal{K}^{(1)}(r;\bm{v})=\frac{\lambda}{\pi}\,\frac{1}{r^{2}+\lambda^{2}}.$ (42)
Equation (42) coincides with the Cauchy kernel [24] on the real line with scale parameter $\lambda$. In particular, the exit law exhibits algebraic decay $\mathcal{K}^{(1)}(r)\sim|r|^{-2}$ and possesses no finite second moment.
A.2 Three-dimensional geometry: absorbing plane
In three spatial dimensions, the absorbing boundary is a plane and the exit law is defined on $\mathbb{R}^{2}$. For constant drift, the boundary kernel derived in Sec. IV admits the representation
$\mathcal{K}^{(2)}(\bm{r};\bm{v})\propto{}$ $\exp\!\left(-\frac{\|\bm{v}\|}{\sigma^{2}}\sqrt{\|\bm{r}\|^{2}+\lambda^{2}}\right)$ (43)
$\times\frac{1+\dfrac{\|\bm{v}\|}{\sigma^{2}}\sqrt{\|\bm{r}\|^{2}+\lambda^{2}}}{\left(\|\bm{r}\|^{2}+\lambda^{2}\right)^{3/2}}.$
In the zero-drift limit, the exponential factor converges to unity and the remaining algebraic kernel dominates. Consequently,
$\lim_{\|\bm{v}\|\to 0}\mathcal{K}^{(2)}(\bm{r};\bm{v})=\frac{\lambda}{2\pi}\frac{1}{\left(\|\bm{r}\|^{2}+\lambda^{2}\right)^{3/2}}.$ (44)
The resulting limit form corresponds to an isotropic two-dimensional Cauchy-type [24] exit law on the boundary plane. As in the two-dimensional case, the kernel exhibits algebraic decay $\mathcal{K}^{(2)}(\bm{r})\sim\|\bm{r}\|^{-3}$ and lacks a finite variance.
Appendix B Planar joint time–location law as a consistency check (drift-free)
Joint vs. marginal.
When a joint density $f_{\tau_{\Omega},\Xi_{\parallel}}(t,y_{\parallel})$ exists, the boundary kernel is recovered by marginalization:
$\mathcal{K}^{(0)}(y_{\parallel})=\int_{0}^{\infty}f_{\tau_{\Omega},\Xi_{\parallel}}(t,y_{\parallel})\,\mathrm{d}t,$
consistent with $\omega^{\bm{x}}(\mathrm{d}y_{\parallel})=\mathcal{K}^{(0)}(y_{\parallel})\,\mathrm{d}y_{\parallel}$.
This appendix provides a simple parabolic consistency check for the generator–Green-function framework developed in the main text. Throughout the paper, the exit time is treated as an internal variable and eliminated at the level of the generator. In the special case of a planar absorbing boundary, however, the drift-free joint law of the exit time and the tangential exit location admits a classical closed-form representation [21, 22]. Marginalizing over time recovers the drift-free (Cauchy-type) exit law on the boundary, consistent with both the planar boundary kernels obtained via the elliptic route in Sec. IV and the asymptotic analysis in Sec. V.
B.1 Drift-free joint law in the half-space
Let $d\geq 2$ denote the ambient dimension in this appendix and consider the absorbing half-space
$\Omega$ $=\{x=(x_{1},x_{\parallel})\in\mathbb{R}\times\mathbb{R}^{d-1}:x_{1}>0\},$ (45a)
$\partial\Omega$ $=\{(0,y_{\parallel}):y_{\parallel}\in\mathbb{R}^{d-1}\}.$ (45b)
The process starts from $\bm{X}_{0}=(\lambda,x_{\parallel})$ with $\lambda>0$. Define the exit time and tangential exit location by
$\tau_{\Omega}:=\inf\{t>0:\bm{X}_{t}\in\partial\Omega\},\quad\Xi_{\parallel}:=(\bm{X}_{\tau_{\Omega}})_{\parallel}\in\mathbb{R}^{d-1}.$ (46)
In the absence of drift, the normal and tangential components of the Brownian motion decouple, so that the exit time is governed solely by the one-dimensional normal motion, while tangential displacements remain unconstrained at the exit event [17]. As a consequence, the joint law of $(\tau_{\Omega},\Xi_{\parallel})$, written in a joint PDF form $f(\cdot,\cdot)$, factorizes as
$f^{0}_{\tau_{\Omega},\Xi_{\parallel}}(t,y_{\parallel})=f^{0}_{\tau_{\Omega}}(t)\,f^{0}_{\Xi_{\parallel}\mid\tau_{\Omega}}(y_{\parallel}\mid t),\quad t>0,\ y_{\parallel}\in\mathbb{R}^{d-1}.$ (47)
Here the one-dimensional first-passage-time density in the normal direction is
$f^{0}_{\tau_{\Omega}}(t)=\frac{\lambda}{\sqrt{4\pi D\,t^{3}}}\exp\!\left(-\frac{\lambda^{2}}{4Dt}\right),$ (48)
while the conditional tangential distribution is the free heat kernel in $\mathbb{R}^{d-1}$,
$f^{0}_{\Xi_{\parallel}\mid\tau_{\Omega}}(y_{\parallel}\mid t)=\frac{1}{(4\pi Dt)^{\frac{d-1}{2}}}\exp\!\left(-\frac{\|y_{\parallel}-x_{\parallel}\|^{2}}{4Dt}\right).$ (49)
Equation (47) is the planar counterpart of the reflection-principle description of first passage [21, 17]—the exit time is governed by the normal motion, while tangential displacements remain unconstrained at the exit event.
B.2 Marginalization and recovery of the drift-free exit law
Marginalizing (47) over the exit time yields the drift-free exit law on $\partial\Omega$:
$\omega^{0}(\mathrm{d}y_{\parallel})$ $=\mathcal{K}^{(0)}(y_{\parallel})\,\mathrm{d}y_{\parallel},$ (B6a)
$\mathcal{K}^{(0)}(y_{\parallel})$ $=\int_{0}^{\infty}f^{0}_{\tau_{\Omega},\Xi_{\parallel}}(t,y_{\parallel})\,\mathrm{d}t.$ (B6b)
Substituting (48)–(49) and defining $\rho^{2}:=\|y_{\parallel}-x_{\parallel}\|^{2}+\lambda^{2}$, we obtain
$\mathcal{K}^{(0)}(y_{\parallel})=\frac{\lambda}{(4\pi D)^{\frac{d}{2}}}\int_{0}^{\infty}t^{-\left(1+\frac{d}{2}\right)}\exp\!\left(-\frac{\rho^{2}}{4Dt}\right)\mathrm{d}t.$ (51)
Using the standard change of variables $s=\rho^{2}/(4Dt)$ and the definition of the Gamma function yields
$\int_{0}^{\infty}t^{-\left(1+\frac{d}{2}\right)}\exp\!\left(-\frac{\rho^{2}}{4Dt}\right)\mathrm{d}t=\left(\frac{4D}{\rho^{2}}\right)^{\frac{d}{2}}\Gamma\!\left(\frac{d}{2}\right),$ (52)
and therefore
$\mathcal{K}^{(0)}(y_{\parallel})=\frac{\lambda\,\Gamma\!\left(\frac{d}{2}\right)}{\pi^{\frac{d}{2}}}\frac{1}{\left(\|y_{\parallel}-x_{\parallel}\|^{2}+\lambda^{2}\right)^{\frac{d}{2}}}.$ (53)
Equation (53) is precisely the drift-free Cauchy-type exit law in the half-space [21, 7, 9]. It is isotropic along the boundary and exhibits algebraic decay $\mathcal{K}^{(0)}(y_{\parallel})\sim\|y_{\parallel}\|^{-d}$ as $\|y_{\parallel}\|\to\infty$, which is consistent with the general tail form $\mathcal{K}^{(0)}_{d}(\bm{r})\sim\|\bm{r}\|^{-(d+1)}$ stated in the main text, noting that here $y_{\parallel}\in\mathbb{R}^{d-1}$ parametrizes the boundary. For $d=2$ it reduces to the standard Cauchy kernel on $\mathbb{R}$, while for $d=3$ it coincides with the planar Poisson kernel on $\mathbb{R}^{2}$. This recovers, at the parabolic (time-resolved) level, the same drift-free exit law obtained in the main text and Appendix A via the generator–Green-function approach.
Remark (scope of the consistency check).
The purpose of this appendix is purely consistency. It verifies, in the planar drift-free setting, that explicitly reintroducing the time variable and marginalizing it out reproduces the same exit law as the generator–Green-function route. No claim is made that an equally simple joint time–location representation persists beyond planar geometries or in the presence of general drift fields.
Table 2: Summary of notation used in this paper.
Symbol Description
$\Omega\subset\mathbb{R}^{d+1}$ Absorbing domain (half-space)
$\partial\Omega$ Absorbing boundary
$\bm{r}$ Tangential displacement along the boundary
$\lambda$ Normal distance to $\partial\Omega$
$d$ Spatial dimension
$\bm{x}\in\Omega$ Initial position
$\bm{y}\in\partial\Omega$ Exit location
$\mathbf{X}_{t}$ Drift–diffusion process
$\tau_{\Omega}$ First hitting time of $\partial\Omega$
$\Xi:=\mathbf{X}_{\tau_{\Omega}}$ Exit location (random variable)
$\omega^{\bm{x}}(\mathrm{d}\bm{y})$ Exit measure on $\partial\Omega$
$\mathcal{K}(\bm{x},\bm{y})$ Boundary kernel density
$\mathrm{d}S_{\bm{y}}$ Surface measure
$\bm{v}$ Constant drift vector
$\bm{u}$ Dimensionless drift
$D$ Diffusion coefficient
$\sigma$ Noise strength (SDE amplitude)
$K_{\nu}(\cdot)$ Modified Bessel function (second kind)
Appendix C Asymptotic Properties of Modified Bessel Functions
This appendix summarizes the asymptotic behavior of the modified Bessel function of the second kind, $K_{\nu}(z)$. We specifically highlight the orders $\nu=1$ and $\nu=3/2$, which correspond to the boundary exit laws in two- and three-dimensional geometries, respectively. These properties underpin the transition between the diffusion-dominated and drift-regularized regimes discussed in the main text.
C.1 Small-Argument Regime ($z\to 0^{+}$): Algebraic Divergence
In the diffusion-dominated regime or at short transverse distances, the argument $z$ is small. For any fixed order $\nu>0$, the function exhibits an algebraic divergence as $z\to 0^{+}$ [1, Eq. (9.6.9)]:
$K_{\nu}(z)\sim\frac{1}{2}\Gamma(\nu)\left(\frac{2}{z}\right)^{\nu}.$ (54)
For the specific cases analyzed in this work, the leading-order behaviors are:
$K_{1}(z)$ $\sim\frac{1}{z},$ (55a)
$K_{3/2}(z)$ $\sim\sqrt{\frac{\pi}{2}}z^{-3/2}.$ (55b)
These power-law divergences, $O(z^{-1})$ and $O(z^{-3/2})$, are the mathematical origin of the scale-free, heavy-tailed behavior characteristic of purely Brownian transport.
C.2 Large-Argument Regime ($z\to\infty$): Exponential Screening
For large arguments, which arise from significant drift or large transmission distances, $K_{\nu}(z)$ is dominated by exponential decay. For any fixed order $\nu$, the general expansion is given by [1, Eq. (9.7.2)]:
$K_{\nu}(z)\sim\sqrt{\frac{\pi}{2z}}\,e^{-z}\left(1+\frac{4\nu^{2}-1}{8z}+\dots\right).$ (56)
Of particular relevance is the case $\nu=3/2$ (3D planar geometry), where the function admits an exact closed-form representation for all $z>0$:
$K_{3/2}(z)=\sqrt{\frac{\pi}{2z}}\,e^{-z}\left(1+\frac{1}{z}\right).$ (57)
From (56) and (57), the leading-order terms as $z\to\infty$ reduce to:
$K_{1}(z)$ $\sim\sqrt{\frac{\pi}{2z}}\,e^{-z},$ (58a)
$K_{3/2}(z)$ $\sim\sqrt{\frac{\pi}{2z}}\,e^{-z}.$ (58b)
The exponential factor $e^{-z}$ represents the drift-induced screening mechanism that regularizes the exit law. In the 3D planar case, the subleading $(1+z^{-1})$ factor in (57) gives rise to the polynomial correction $(1+\|\bm{u}\|\rho)$ appearing in the boundary kernel, cf. Eq. (30).

### v2

@@
 In diffusion-dominated regimes, the exit law exhibits scale-free, heavy-tailed spatial fluctuations along the boundary, whereas a nonzero drift component introduces an intrinsic length scale that suppresses these tails and reorganizes the exit statistics.
-Within a generator-based formulation, the FHL arises naturally as a boundary measure induced by an elliptic operator, allowing explicit boundary kernels to be derived analytically in canonical geometries.
-Planar absorbing boundaries in different ambient dimensions are examined as benchmark cases, illustrating how directed transport regularizes diffusion-driven fluctuations and induces qualitative transitions in boundary statistics.
+Within a generator-based formulation, the FHL arises naturally as a boundary measure induced by an elliptic operator, allowing exact $(d+1)$-dimensional boundary kernels to be derived analytically.
+Planar absorbing boundaries are examined as benchmark cases and validated via Monte Carlo simulations, illustrating how directed transport thermodynamically regularizes diffusion-driven fluctuations—quantified by a robust effective width—and induces qualitative transitions in boundary statistics.
 Overall, the present work provides a unified structural framework for first-hitting location laws and highlights FHL statistics as natural probes of geometry, drift, and irreversibility in stochastic transport.
@@
 In particular, the planar cases make transparent how directed transport suppresses diffusion-induced spatial fluctuations along the boundary and introduces a characteristic length scale $\ell_{u}=\sigma^{2}/\|\bm{v}\|=1/u$, which separates diffusion-dominated, scale-free behavior from drift-regularized regimes with exponentially screened tails.
+Novelty and main contributions.
+Because the Green-function and exit-measure representations are closely related to classical potential-theory concepts, it is important to clarify what is newly synthesized here for the stochastic transport and statistical physics communities.
+The primary advance of this work lies in systematically treating the FHL itself as the primary physical observable, utilizing a rigorous stochastic framework that goes beyond macroscopic transport equations.
+Specifically, we highlight the following contributions:
+  • Necessity of the stochastic framework: We provide a unified derivation that cleanly separates the spatial boundary measure from temporal stopping-time problems.
+We demonstrate that deriving the exact physical exit law requires a crucial exponential tilting conversion (mathematically equivalent to the Girsanov theorem [21]).
+This highlights that macroscopic deterministic PDEs alone are insufficient to naturally capture this measure, justifying our generator-based stochastic approach.
+  • General $(d+1)$-dimensional closed-form kernels: We derive explicit, analytically closed-form half-space Poisson kernels for constant drift in arbitrary $(d+1)$ dimensions.
+This result fills a documented gap in foundational mathematical physics literature, which has primarily restricted explicit analytical evaluations to low dimensions ($d\leq 3$), see [23].
+  • Thermodynamic regularization and numerical validation: We identify a drift-induced crossover between scale-free diffusion and exponentially screened transport regimes.
+We explicitly demonstrate how directed transport acts as a localized probe, physically compressing scale-free Cauchy heavy tails into an exponentially screened spatial footprint.
+These analytical phenomena are explicitly validated by particle-based Monte Carlo simulations of the underlying Langevin dynamics.
+  • Robust geometric diagnostics: We introduce entropy-based diagnostics built on the exit-location distribution.
+By defining the “effective width” ($W_{\text{eff}}=\exp(H)$), we provide a robust, finite geometric measure of spatial dispersion that remains well-behaved even in the Cauchy limit, effectively resolving the divergence issues associated with traditional variance-based metrics.
+A concise summary of these methodological shifts compared to the standard parabolic PDE route is provided in Table 2.
+Furthermore, to bridge the terminology gap between rigorous stochastic analysis and macroscopic physical intuition, Table 3 provides a methodological dictionary translating our stochastic formalism into standard transport concepts.
 Scope and limitations.
@@
 While the generator-based formulation and the qualitative transition between diffusion-dominated and drift-regularized regimes are not restricted to planar domains, explicit closed-form boundary kernels are derived here only for the half-space geometry.
-Extensions to curved boundaries, time-dependent drift fields, and fully information-theoretic performance measures such as channel capacity are beyond the scope of this study and will be addressed elsewhere.
+Extensions to curved boundaries, time-dependent drift fields, and fully information-theoretic performance measures such as channel capacity are beyond the scope of this study and will be addressed in future work.
 Our aim is instead to isolate the minimal structural mechanisms by which geometry and drift jointly shape boundary exit laws and their induced observables.
+Paper organization.
 The remainder of the paper is organized as follows.
@@
 VI concludes the paper.
-Table 1: Parabolic PDE versus generator-based routes to first-hitting location laws
-Parabolic PDE route Generator–Green-function route
-Step 1 Construct a free-space fundamental solution for a parabolic partial differential equation (PDE). Pass to the infinitesimal generator of the Markov semigroup.
-Step 2 Impose absorbing boundary conditions (e.g., by image methods). Solve an elliptic boundary-value problem with Dirichlet data.
-Step 3 Compute boundary flux and integrate over all time. Extract the boundary kernel from the normal derivative of the associated Green function.
 II Model
@@
 The boundary is therefore a (d1)(d-1)-dimensional manifold, and boundary-parallel coordinates are denoted by $\bm{r}\in\mathbb{R}^{d-1}$.
-A summary of notation is provided in Table II for reference.
+A summary of notation is provided in Table 1 for reference.
 The diffusion coefficient is denoted by $D>0$, with the standard parametrization $D=\sigma^{2}/2$, following the conventional usage in statistical physics whereby $D$ characterizes the macroscopic spreading rate of the process and appears as the prefactor of the Laplacian in the diffusion equation.
@@
 With this convention, the cases $d=2$ and $d=3$ correspond respectively to an absorbing line and an absorbing plane, while higher-dimensional extensions fit within the same unified framework.
+Table 1: Summary of key notation.
+Symbol Description
+$\Omega\subset\mathbb{R}^{d+1}$ Absorbing domain (half-space)
+$\partial\Omega$ Absorbing boundary
+$\bm{r}$ Tangential displacement along the boundary
+$\lambda$ Normal distance to $\partial\Omega$
+$d$ Spatial dimension
+$\bm{x}\in\Omega$ Initial position
+$\bm{y}\in\partial\Omega$ Exit location
+$\mathbf{X}_{t}$ Drift–diffusion process
+$\tau_{\Omega}$ First hitting time of $\partial\Omega$
+$\Xi:=\mathbf{X}_{\tau_{\Omega}}$ Exit location (random variable)
+$\omega^{\bm{x}}(\mathrm{d}\bm{y})$ Exit measure on $\partial\Omega$
+$\mathcal{K}(\bm{x},\bm{y})$ Boundary kernel density
+$\mathrm{d}S_{\bm{y}}$ Surface measure
+$\bm{v}$ Constant drift vector
+$\bm{u}$ Dimensionless drift
+$D$ Diffusion coefficient
+$\sigma$ Noise strength (SDE amplitude)
+$K_{\nu}(\cdot)$ Modified Bessel function (second kind)
 II.1 Stochastic transport model
@@
 III A unified framework for determining first-hitting location laws
-This section presents a unified route, summarized in Table I, for determining FHL laws induced by drift–diffusion processes in domains with absorbing boundaries.
+Table 2: Parabolic PDE versus generator-based routes to first-hitting location laws
+Parabolic PDE route Generator–Green-function route
+Step 1 Construct a free-space fundamental solution for a parabolic partial differential equation (PDE). Pass to the infinitesimal generator of the Markov semigroup.
+Step 2 Impose absorbing boundary conditions (e.g., by image methods). Solve an elliptic boundary-value problem with Dirichlet data.
+Step 3 Compute boundary flux and integrate over all time. Extract the boundary kernel from the normal derivative of the associated Green function.
+This section presents a unified route, summarized in Table 2, for determining FHL laws induced by drift–diffusion processes in domains with absorbing boundaries.
 Related generator-based descriptions of boundary observables and boundary local time have been extensively studied in the context of reflected and absorbed diffusions; see, e.g., [12].
+To facilitate this discussion for the broad statistical physics community, Table 3 provides a direct mapping between the stochastic operators used in our framework and their macroscopic transport equivalents.
 We refer to the exit location on the boundary as the FHL.
@@
 Drift introduces an intrinsic length scale and exponentially suppresses large boundary displacements, leading to localized and anisotropic hitting profiles consistent with the drift-regularized analytical kernels.
-IV.4 General $(d+1)$-dimensional structure
+IV.4 General $(d+1)$-dimensional exit law
 The explicit results in two and three dimensions reveal a common, dimension-independent structure of boundary exit laws in the absorbing half-space.
 In both cases, the ambient space has dimension $d+1$, while the induced boundary kernel lives on a $d$-dimensional absorbing interface.
-Guided by these expressions, we are naturally led to the conjecture that in $\mathbb{R}^{d+1}$ the boundary kernel admits the universal form
+Guided by low dimensional expressions, we are naturally led to the following general expression for the boundary kernel in $\mathbb{R}^{d+1}$:
 $\mathcal{K}^{(d)}(\bm{r};\bm{u},\lambda)$ $=2\lambda\frac{\|\bm{u}\|^{\frac{d+1}{2}}}{(2\pi)^{\frac{d+1}{2}}}\exp\!\big(\bm{u}_{\parallel}^{\mathsf{T}}\bm{r}-u_{d}\lambda\big)$
@@
 Here $\bm{r}\in\mathbb{R}^{d}$ denotes the tangential displacement along the absorbing boundary.
-While (IV.4) is stated as a conjecture, its structure is strongly constrained by symmetry and scaling considerations.
+A concise derivation of Eq.
+(IV.4) based on the heat-kernel resolvent representation of the modified Helmholtz operator is provided in Appendix D.
+The structure of the kernel is strongly constrained by symmetry and scaling considerations.
 In particular, translational invariance along the absorbing boundary, rotational symmetry in the $d$ boundary-parallel directions, and the exponential tilting induced by the drift together single out the modified Bessel kernel appearing above.
-From this perspective, (IV.4) should be viewed as the natural extension of the exact two- and three-dimensional results to general boundary dimension, rather than as a heuristic guess.
-As a sanity check, we specialize the conjectured expression (IV.4) to the case $d=1$, corresponding to a two-dimensional ambient space with a one-dimensional absorbing boundary.
+From this perspective, Eq.
+(IV.4) should be viewed as the natural extension of the exact two- and three-dimensional results to general boundary dimension.
+As a sanity check, we specialize the general expression (IV.4) to the case $d=1$, corresponding to a two-dimensional ambient space with a one-dimensional absorbing boundary.
 In this setting, the tangential displacement reduces to a scalar $r\in\mathbb{R}$, and the modified Bessel function appearing in (IV.4) becomes $K_{1}$ [1, 22].
@@
 (30).
-Together with the $d=1$ case, this agreement confirms that the conjectured form (IV.4) correctly reproduces the known closed-form results in both two and three dimensions.
+Together with the $d=1$ case, this agreement confirms that Eq.
+(IV.4) correctly reproduces the known closed-form results in both two and three dimensions.
+Figure 5: Numerical validation of the FHL distribution on a three-dimensional planar absorbing boundary.
+(a) Empirical PDF obtained from $10^{6}$ Monte Carlo trajectories of the Langevin dynamics Eq.
+(2).
+(b) Analytical boundary kernel calculated via Eq.
+(32).
+The parameters are set to $\lambda=1.0\,\mu\text{m}$, $\sigma=1.0$, and a constant drift $\mathbf{v}=[2,-3,1]\,\mu\text{m/s}$.
+The simulation results are in excellent agreement with the analytical prediction, accurately capturing the drift-induced anisotropy and the localized hitting profile.
+IV.5 Numerical validation via Monte Carlo simulation
+To validate the analytical boundary kernels derived in this section, we perform particle-based Monte Carlo simulations of the underlying Langevin dynamics Eq.
+(2).
+We consider a 3D system with an absorbing plane at $x_{3}=\lambda=1.0\,\mu\text{m}$ and a noise strength $\sigma=1.0$.
+A uniform drift $\mathbf{v}=[2,-3,1]\,\mu\text{m/s}$ is applied to introduce both normal and tangential transport components.
+As shown in Fig.
+5, the spatial distribution of the first-hitting locations obtained from $10^{6}$ independent trajectories is in excellent agreement with the analytical prediction of Eq.
+(30).
+Specifically, Fig.
+5(a) displays the empirical probability density function (PDF) sampled from the simulated hitting events, while Fig.
+5(b) shows the corresponding theoretical kernel.
+In particular, the theory accurately captures the drift-induced anisotropy and the exponential suppression of tangential excursions along the planar interface.
+This agreement confirms that the generator-based framework provides a robust structural description of the microscopic stochastic transport process, effectively bridging the path-wise Langevin dynamics with the macroscopic boundary exit law.
 V Asymptotic structure and information observables of exit laws
@@
 From a broader statistical-physics perspective, these observables may be viewed as geometric manifestations of nonequilibrium first-passage dynamics, rather than as quantities tied to externally imposed communication protocols [2, 10, 24, 5].
+Figure 6: Spatial localization and effective width compression.
+(a) Exit density $K^{(1)}(r)$ for varying drift strengths $u$, showing the transition from scale-free Cauchy decay ($u=0$) to exponential screening.
+(b) Effective width $W_{\text{eff}}=\exp(H)$ as a function of drift strength $u$.
+The sharp decrease quantifies the drift-induced regularization of boundary fluctuations.
+Unlike the variance, which diverges in the Cauchy limit (red circle), $W_{\text{eff}}$ provides a robust and finite geometric measure of the spatial spread across all transport regimes.
 V.1 Asymptotic structure of planar exit laws
@@
 In this sense, $\ell_{u}$ acts as a crossover length separating diffusion-dominated and drift-regularized behavior at the level of boundary exit laws.
-V.2 Differential entropy of the exit location
-Let $\Xi$ denote the boundary exit location,
-$\Xi:=\mathbf{X}_{\tau_{\Omega}}\in\partial\Omega,\quad\tau_{\Omega}:=\inf\{t>0:\mathbf{X}_{t}\in\partial\Omega\},$ (38)
-under an initial condition $\mathbf{X}_{0}=\bm{x}$.
-Whenever the exit law $\omega^{\bm{x}}$ is absolutely continuous with respect to the surface measure on $\partial\Omega$, we write $\omega^{\bm{x}}(\mathrm{d}\bm{y})=\mathcal{K}(\bm{x},\bm{y})\mathrm{d}S_{\bm{y}}$ and define the (differential) entropy [6] of the exit location as
+V.2 Differential entropy and effective width of the exit location
+To characterize the intrinsic spatial uncertainty of boundary exit events, we treat the first-hitting location $\Xi:=\mathbf{X}_{\tau_{\Omega}}\in\partial\Omega$ as a primary observable.
+Under an initial condition $\mathbf{X}_{0}=\bm{x}$, the exit law $\omega^{\bm{x}}$ is absolutely continuous, allowing us to define the (differential) entropy [7] of the exit location as:
 $H(\Xi\,|\,\bm{x})=-\int_{\partial\Omega}\mathcal{K}(\bm{x},\bm{y})\log\mathcal{K}(\bm{x},\bm{y})\,\mathrm{d}S_{\bm{y}}.$ (38)
-This quantity characterizes the intrinsic spatial uncertainty of the exit event along the absorbing boundary and serves as a natural information observable induced by the boundary-hitting statistics.
-In diffusion-dominated regimes, the exit law exhibits an algebraically decaying heavy tail along the boundary.
-In the drift-free limit, the kernel reduces to a Cauchy-type form with power-law decay, as in (34).
-Despite the absence of finite moments of order one or higher, such heavy-tailed distributions possess a well-defined differential entropy that depends only logarithmically on the scale parameter.
-The pathological feature relevant to information transmission is therefore not entropy divergence, but rather the lack of an intrinsic spatial scale and the resulting dominance of rare but arbitrarily large tangential displacements.
-In this regime, moment-based characterizations and Gaussian approximations break down, and information measures become highly sensitive to tail behavior.
-The introduction of drift fundamentally alters this picture.
-As demonstrated by the explicit planar kernels in Sec.
-IV, a nonzero drift component directed toward the absorbing boundary induces an exponential suppression of large boundary displacements.
-This screening effect suggests a characteristic length scale $\ell_{u}$, as defined in (37), which governs the effective spatial extent of the exit distribution.
-For $\|\bm{v}\|>0$, the exit law is no longer scale-free, and the spatial uncertainty of the exit location is controlled by $\ell_{u}$.
-In particular, $\ell_{u}$ sets the effective spatial support over which the exit density contributes appreciably to entropy-like functionals, so that increasing drift strength progressively localizes the boundary distribution and reduces the contribution of rare, large displacements.
-This regularization mechanism is already apparent at the level of the boundary kernel itself: the emergence of an intrinsic length scale suppresses long tangential excursions and converts an intrinsically scale-free exit observable into one governed by controlled, length-scale-dependent statistics.
-A fully quantitative analysis of entropy variation or input–output information measures (e.g., the Shannon capacity [6]) would require specifying an input ensemble, and therefore lies beyond the scope of the present work.
+Unlike temporal first-passage metrics, this information observable quantifies the spatial dispersion shaped by the interplay of directed transport and domain geometry.
+In diffusion-dominated regimes, the exit law exhibits an algebraically decaying heavy tail.
+In the zero-drift limit ($\bm{u}=0$), the kernel reduces to a scale-free Cauchy-type form, see Eq.
+(34).
+A critical physical observation is that for such heavy-tailed distributions, traditional moment-based measures of spread (e.g., variance) either diverge or become highly sensitive to rare, long-range excursions.
+Differential entropy, by contrast, depends only logarithmically on the scale parameter and remains finite and well-behaved even in the Cauchy limit, offering a more robust diagnostic.
+To make this measure physically concrete, we relate the entropy to a geometric length scale by defining the effective width (or effective occupied area) of the exit distribution:
+$W_{\text{eff}}=\exp[H(\Xi\,|\,\bm{x})].$ (39)
+In information theory, $W_{\text{eff}}$ corresponds to the volume of the typical set [7].
+Physically, it characterizes the spatial footprint of the hitting pattern that contributes most to the probability mass—intuitively, the effective number of occupied bins at a fixed resolution.
+The emergence of the intrinsic length scale $\ell_{u}=1/u$ fundamentally regularizes this spatial footprint.
+As illustrated in Fig.
+6(a) for a one-dimensional absorbing boundary with a fixed initial normal distance $\lambda=1$, increasing the drift $u$ toward the boundary induces exponential tail screening, effectively “compressing” the distribution.
+To ensure numerical reproducibility, $W_{\text{eff}}$ is evaluated by numerically integrating the exact analytical kernel $\mathcal{K}^{(1)}(r;u,\lambda)$ over the infinite domain $(-\infty,\infty)$ using adaptive quadrature routines.
+As shown in Fig.
+6(b), the computed $W_{\text{eff}}$ exhibits a sharp, monotonic decrease from its maximum in the scale-free Cauchy limit.
+This confirms that directed transport acts as a localized probe, suppressing diffusion-driven fluctuations and confining exit events to a region governed by $\ell_{u}$.
+In summary, $W_{\text{eff}}$ serves as a robust geometric summary of the transition from scale-free to drift-dominated transport.
+While it characterizes the intrinsic spread of a single point source, a fully quantitative analysis of input–output information measures, such as the Shannon capacity [7], requires specifying an input ensemble and remains beyond the structural scope of this work.
 VI Conclusion and discussion
 This work examined boundary exit laws associated with the location of first-passage events in drift–diffusion processes on extended domains, with particular emphasis on their asymptotic structure and induced information observables.
-Within a generator–Green-function framework, we derived explicit boundary kernels for planar absorbing geometries and characterized the resulting exit measures across diffusion-dominated and drift-regularized regimes.
+Within a generator–Green-function framework, we derived exact $(d+1)$-dimensional boundary kernels for planar absorbing geometries and validated them via Monte Carlo simulations, characterizing the resulting exit measures across diffusion-dominated and drift-regularized regimes.
 Viewed from this perspective, the exit location, rather than the exit time alone, emerges as a sensitive observable through which geometry and directed transport are imprinted on boundary statistics.
@@
 In this sense, first-hitting location statistics provide a concrete boundary manifestation of irreversibility in stochastic transport, reflecting how drift reorganizes spatial fluctuations at absorbing interfaces.
-Information-theoretic quantities, such as the differential entropy of the exit location, serve as convenient diagnostics of this transition, reflecting changes in boundary statistics induced jointly by stochastic transport and geometry.
+Information-theoretic quantities, such as the differential entropy of the exit location and the associated effective width, serve as robust geometric diagnostics of this thermodynamic regularization.
+They effectively capture changes in boundary statistics induced jointly by stochastic transport and geometry, remaining well-behaved even when traditional variance-based metrics diverge.
 The planar results presented here provide benchmark solutions that capture the leading-order structure of boundary exit laws in more general geometries, with curvature and global shape effects entering as higher-order corrections.
 More broadly, the framework developed in this work identifies first-hitting location statistics as a natural class of boundary-induced observables through which geometric and dynamical aspects of stochastic transport can be examined within a unified setting.
-Extensions to curved absorbing boundaries and to time-dependent drift constitute natural directions for future work, where additional geometric and dynamical scales are expected to further enrich the structure of boundary exit laws.
+From a practical standpoint, the numerical evaluation of these boundary kernels in arbitrary geometries connects to recent computational research [4] on modified Helmholtz operators.
+Extensions to curved absorbing boundaries, the implementation of boundary integral methods for general domains, and to time-dependent drift constitute natural directions for future work, where additional geometric and dynamical scales are expected to further enrich the structure of boundary exit laws.
 Acknowledgements.
@@
 Throughout the main text, the exit law on $\partial\Omega$ is written as a boundary measure $\omega^{\bm{x}}(\mathrm{d}\bm{y})=\mathcal{K}(\bm{x},\bm{y})\,\mathrm{d}S_{\bm{y}}$.
-In the appendices we sometimes parameterize $\partial\Omega$ by local coordinates $y\in\mathbb{R}^{d-1}$ (e.g., $y=r$ for $d=2$ and $y=\bm{r}$ for $d=3$), in which case $\mathrm{d}S_{\bm{y}}=\mathrm{d}y$ and the same law is written in probability density function (PDF) form $f_{\Xi}(y)=\mathcal{K}(\bm{x},\bm{y}(y))$.
+In the appendices we sometimes parameterize $\partial\Omega$ by local coordinates $y\in\mathbb{R}^{d-1}$ (e.g., $y=r$ for $d=2$ and $y=\bm{r}$ for $d=3$), in which case $\mathrm{d}S_{\bm{y}}=\mathrm{d}y$ and the same law is written in PDF form $f_{\Xi}(y)=\mathcal{K}(\bm{x},\bm{y}(y))$.
 More generally, for a coordinate map $y\mapsto\bm{y}(y)$, the density takes the form $f_{\Xi}(y)=\mathcal{K}(\bm{x},\bm{y}(y))\,J(y)$, where $J(y)$ denotes the surface Jacobian.
@@
 As in the two-dimensional case, the kernel exhibits algebraic decay $\mathcal{K}^{(2)}(\bm{r})\sim\|\bm{r}\|^{-3}$ and lacks a finite variance.
-Appendix B Planar joint time–location law as a consistency check (drift-free)
+Appendix B Planar Joint Time–Location Law as a Consistency Check (Drift-Free Case)
 Joint vs. marginal.
@@
 No claim is made that an equally simple joint time–location representation persists beyond planar geometries or in the presence of general drift fields.
-Table 2: Summary of notation used in this paper.
-Symbol Description
-$\Omega\subset\mathbb{R}^{d+1}$ Absorbing domain (half-space)
-$\partial\Omega$ Absorbing boundary
-$\bm{r}$ Tangential displacement along the boundary
-$\lambda$ Normal distance to $\partial\Omega$
-$d$ Spatial dimension
-$\bm{x}\in\Omega$ Initial position
-$\bm{y}\in\partial\Omega$ Exit location
-$\mathbf{X}_{t}$ Drift–diffusion process
-$\tau_{\Omega}$ First hitting time of $\partial\Omega$
-$\Xi:=\mathbf{X}_{\tau_{\Omega}}$ Exit location (random variable)
-$\omega^{\bm{x}}(\mathrm{d}\bm{y})$ Exit measure on $\partial\Omega$
-$\mathcal{K}(\bm{x},\bm{y})$ Boundary kernel density
-$\mathrm{d}S_{\bm{y}}$ Surface measure
-$\bm{v}$ Constant drift vector
-$\bm{u}$ Dimensionless drift
-$D$ Diffusion coefficient
-$\sigma$ Noise strength (SDE amplitude)
-$K_{\nu}(\cdot)$ Modified Bessel function (second kind)
 Appendix C Asymptotic Properties of Modified Bessel Functions
@@
 (30).
+Table 3: Methodological dictionary connecting the stochastic-process formulation to macroscopic transport concepts.
+This mapping translates the trajectory-level probabilistic operators and boundary measures used in this work into the ensemble-level flux and transport language familiar to the statistical physics community.
+Stochastic Formulation Macroscopic Transport Equivalent Physical Interpretation
+Infinitesimal generator $\mathcal{L}$ Steady-state advection–diffusion operator Local bulk transport dynamics (drift and diffusion) governing the system.
+Transition semigroup $P_{t}$ Time-evolution (Green’s function) operator Time progression of the ensemble-averaged probability density function.
+Dirichlet Green function $\mathcal{G}(\mathbf{x},\mathbf{y})$ Steady-state concentration field Bulk spatial distribution maintained by a continuous point source at $\mathbf{x}$.
+Dynkin’s formula Global flux-balance identity Conservation law equating bulk transport evolution to the total boundary flux.
+Exit measure $\omega^{\mathbf{x}}(\mathrm{d}\mathbf{y})$ and Kernel $\mathcal{K}(\mathbf{x},\mathbf{y})$ Normal boundary flux distribution Spatial profile of the steady-state absorption current across the interface.
+First-hitting location $\mathbf{X}_{\tau_{\Omega}}$ Point of absorption / Exit coordinate The path-wise, trajectory-level spatial observable of the exit event.
+Appendix D Derivation of the General-Dimension Exit Kernel via the Heat-Kernel Resolvent
+This appendix provides a rigorous and self-contained derivation of the $(d+1)$-dimensional explicit boundary kernel presented in Eq.
+(IV.4).
+Instead of postulating the free-space Green’s function for the modified Helmholtz operator, we derive the boundary kernel by evaluating the resolvent of the purely diffusive heat kernel, an approach that is standard in stochastic transport.
+Let $d+1$ denote the ambient spatial dimension.
+We consider the fundamental solution to the diffusion equation corresponding to the generator $\frac{\sigma^{2}}{2}\Delta$ in $\mathbb{R}^{d+1}$, which represents the transition density of a Brownian motion with diffusion coefficient $D=\sigma^{2}/2$ [20, 9].
+Namely,
+$p(t,\bm{x},\bm{y})=\frac{1}{(2\pi\sigma^{2}t)^{\frac{d+1}{2}}}\exp\!\left(-\frac{\|\bm{x}-\bm{y}\|^{2}}{2\sigma^{2}t}\right).$ (59)
+To satisfy the absorbing (Dirichlet) boundary condition on the boundary $\partial\Omega$, we use the method of images [14].
+By placing an image sink at $\bm{x}^{*}$, symmetrically opposite to the initial position $\bm{x}$ with respect to the boundary, the transition density of the absorbed process is
+$p_{\text{dir}}(t,\bm{x},\bm{y})=p(t,\bm{x},\bm{y})-p(t,\bm{x}^{*},\bm{y}).$ (60)
+For a boundary point $\bm{y}\in\partial\Omega$, let $\bm{r}\in\mathbb{R}^{d}$ denote the tangential displacement and let $\lambda>0$ denote the initial normal distance.
+The distance from both the source and the image to the boundary point is $\rho=\sqrt{\|\bm{r}\|^{2}+\lambda^{2}}$.
+Evaluating the outward normal derivative $\partial_{n(\bm{y})}$ of $p_{\text{dir}}$ at the boundary yields
+$\partial_{n(\bm{y})}p_{\text{dir}}(t,\bm{x},\bm{y})=\frac{2\lambda}{\sigma^{2}t}\frac{1}{(2\pi\sigma^{2}t)^{\frac{d+1}{2}}}\exp\!\left(-\frac{\rho^{2}}{2\sigma^{2}t}\right).$ (61)
+To connect this to the elliptic boundary-value problem formulated in Eq.
+(22), we construct the Green function $\tilde{\mathcal{G}}(\bm{x},\bm{y})$ satisfying $(\Delta-\|\bm{u}\|^{2})\tilde{\mathcal{G}}=-\delta(\bm{x}-\bm{y})$.
+Using the resolvent of the diffusion generator, this Green function is given by the time integral of the absorbed heat kernel discounted by the rate $\alpha=\frac{\sigma^{2}}{2}\|\bm{u}\|^{2}$, and scaled by $\frac{\sigma^{2}}{2}$ to match the operator definition,
+$\tilde{\mathcal{G}}(\bm{x},\bm{y})=\int_{0}^{\infty}\frac{\sigma^{2}}{2}e^{-\frac{\sigma^{2}}{2}\|\bm{u}\|^{2}t}p_{\text{dir}}(t,\bm{x},\bm{y})\,dt.$ (62)
+Consequently, its outward normal derivative is
+$-\partial_{n(\bm{y})}\tilde{\mathcal{G}}(\bm{x},\bm{y})$ (63)
+$\quad=\int_{0}^{\infty}\frac{\sigma^{2}}{2}e^{-\frac{\sigma^{2}}{2}\|\bm{u}\|^{2}t}\bigl[-\partial_{n(\bm{y})}p_{\text{dir}}(t,\bm{x},\bm{y})\bigr]\,dt$
+$\quad=\frac{\lambda}{(2\pi\sigma^{2})^{\frac{d+1}{2}}}\int_{0}^{\infty}t^{-\frac{d+3}{2}}$
+$\qquad\times\exp\!\left(-\frac{\sigma^{2}}{2}\|\bm{u}\|^{2}t-\frac{\rho^{2}}{2\sigma^{2}t}\right)dt.$
+This integral can be evaluated analytically using the standard integral representation of the modified Bessel function of the second kind [1],
+$\int_{0}^{\infty}\tau^{-\nu-1}\exp\!\left(-p\tau-\frac{q}{\tau}\right)d\tau=2\left(\frac{p}{q}\right)^{\nu/2}K_{\nu}\!\left(2\sqrt{pq}\right).$ (64)
+Substituting $\nu=\frac{d+1}{2}$, $p=\frac{\sigma^{2}}{2}\|\bm{u}\|^{2}$, and $q=\frac{\rho^{2}}{2\sigma^{2}}$ into Eq.
+(63), the argument of the Bessel function evaluates to $2\sqrt{pq}=\|\bm{u}\|\rho$.
+Remarkably, the prefactor evaluates to $(p/q)^{\nu/2}=(\sigma^{2}\|\bm{u}\|/\rho)^{\frac{d+1}{2}}$, perfectly canceling the explicit $\sigma^{2}$ dependence in the denominator.
+This yields the scale-regularized boundary Poisson kernel:
+$\partial_{n(\bm{y})}\tilde{\mathcal{G}}(\bm{x},\bm{y})=2\lambda\frac{\|\bm{u}\|^{\frac{d+1}{2}}}{(2\pi)^{\frac{d+1}{2}}}\frac{K_{\frac{d+1}{2}}\!\left(\|\bm{u}\|\rho\right)}{\rho^{\frac{d+1}{2}}}.$ (65)
+Finally, to recover the physical first-hitting location boundary kernel, we apply the exponential tilting relation established in Eq.
+(23).
+The displacement vector $\bm{y}-\bm{x}$ corresponds to a tangential shift $\bm{r}$ and a normal shift $\lambda$.
+Multiplying Eq.
+(65) by the exponential factor $\exp(\bm{u}^{\mathsf{\scriptstyle{T}}}(\bm{y}-\bm{x}))=\exp(\bm{u}_{\parallel}^{\mathsf{\scriptstyle{T}}}\bm{r}-u_{d}\lambda)$ yields
+$\mathcal{K}^{(d)}(\bm{r};\bm{u},\lambda)$ $=2\lambda\frac{\|\bm{u}\|^{\frac{d+1}{2}}}{(2\pi)^{\frac{d+1}{2}}}\exp\!\big(\bm{u}_{\parallel}^{\mathsf{\scriptstyle{T}}}\bm{r}-u_{d}\lambda\big)$ (66)
+$\quad\times\frac{K_{\frac{d+1}{2}}\!\left(\|\bm{u}\|\rho\right)}{\rho^{\frac{d+1}{2}}}.$
+This coincides exactly with the closed-form expression in Eq.
+(IV.4).
+The derivation rigorously demonstrates that the high-dimensional exit law naturally resolves into a unified modified Bessel kernel structure, originating fundamentally from the integration over stochastic transport times.